Bond Discount, Entries for Bonds Payable Transactions

On July 1, 20Y1, Livingston Corporation, a wholesaler of manufacturing equipment, issued $4,400,000 of 8-year, 11% bonds at a market (effective) interest rate of 12%, receiving cash of $4,177,688. Interest on the bonds is payable semiannually on December 31 and June 30. The fiscal year of the company is the calendar year.

Required:

For all journal entries, if an amount box does not require an entry, leave it blank.

1. Journalize the entry to record the amount of cash proceeds from the issuance of the bonds on July 1, 20Y1.

Cash   
Discount on Bonds Payable
Bonds Payable

2. Journalize the entries to record the following:

a. The first semiannual interest payment on December 31, 20Y1, and the amortization of the bond discount, using the straight-line method. Round to the nearest dollar.

Interest Expense
Discount on Bonds Payable
Cash

b. The interest payment on June 30, 20Y2, and the amortization of the bond discount, using the straight-line method. Round to the nearest dollar.

Interest Expense
Discount on Bonds Payable
Cash

3. Determine the total interest expense for 20Y1. Round to the nearest dollar.
$

4. Will the bond proceeds always be less than the face amount of the bonds when the contract rate is less than the market rate of interest?
Yes

5. Compute the price of $4,177,688 received for the bonds by using the present value tables in Appendix A. Round your PV values to 5 decimal places and the final answers to the nearest dollar. Your total may vary slightly from the price given due to rounding differences.

Present value of the face amount $
Present value of the semiannual interest payments
Price received for the bonds $
Answer #1
1) Date Account titles & Explanations Debit Credit
1-Jul Cash 4,177,688
Discount on bonds 222,312
bonds payable 4,400,000
2) 31-Dec interest expense 255895
discount on bonds (222312/16) 13895
cash (4,400,000*5.5%) 242000
30-Jun interest expense 255895
discount on bonds 13895
cash 242000
3) total interest expense for 20Y1 255895
4) Yes
5)
Principal 4,400,000
interest 242000
Calculation of bond issue price
Where
i= 6.00%
t= 16
principal * PV of $1 at 6% for 16 yrs =
4,400,000 * 0.39365        = 1732060
interest * PV of ordinary annuity at 6%=
242000 * 10.1059 = 2445628
bond issue price 4177688
so
present value of the face amount 1732060
present value of the semi-annual interest payments 2445628
price received for bonds 4177688 answer